OpenStudy (anonymous):

Find a polynomial p of degree 3 so that p(5) = 153, p(−5) = −307, p(3) = 13, p(−1) = −3, p(x)= use the polynomial to find p(4) p(4)=

OpenStudy (anonymous):

do we know if the polynomial is quadratic or binomial or what type of polynomial P would be?

OpenStudy (anonymous):

@DemolisionWolf I am not quite sure, but it says here that "You can find p by setting p(x) = ax3+bx2+cx+d, and substituting the given values of x and p(x) to obtain 4 equations in the 4 variables a...d."

OpenStudy (anonymous):

p(x) = ax3+bx2+cx+d p(5) = 5^3+5^2+5 = 155, so d would become -2 thus p(5) = 5^3+5^2+5 -2 = 153

OpenStudy (anonymous):

holy crap, is that so? All this time I would assume we put it into a matrix obtained from r+rx+rx^2

OpenStudy (anonymous):

but how would you find p(x) @DemolisionWolf .

OpenStudy (anonymous):

isn't p(x) given to you as p(x) = ax3+bx2+cx+d ?

OpenStudy (anonymous):

yes, but we have to find p(x) from those 4 equations given

OpenStudy (anonymous):

so we need to find a, b, c, and d using a matrix then?

OpenStudy (anonymous):

yeah i guess so

OpenStudy (anonymous):

p(x) = ax3+bx2+cx+d p(5) = A*5^3 + B*5^2 + C*5 + D = 153 p(-5) = A*5^3 + B*5^2 + C*5 + D = -307 p(3) = A*3^3 + B*3^2 + C*3 + D = 13 p(−1) = A*(-1)^3 + B*(-1)^2 + C*(-1) + D= −3 4 equations, 4 unknowns, so the 'x' of each term becomes the 'constant' so to speak, and we are needing to find A, B, C, D, so I belive the first row would look like: [ 5^3 5^2 5 1 153 ] [ 5^3 5^2 5 1 -307 ] [ 3^3 3^2 3 1 13 ] [-1^3 -1^2 -1 1 -3 ] get this matrix into identity form, and you can then solve for A, B, C, D

OpenStudy (anonymous):

right on. that is exactly what i thought.

OpenStudy (anonymous):

sorry it took me so long to understand what you were saying

OpenStudy (anonymous):

yeah its no problem at all.